# Position: Beyond Euclidean - Foundation Models Should Embrace Non-Euclidean Geometries

### v1

@@ A.2 Non-Applicability of the Nash Embedding Theorem
The Nash Embedding Theorem roughly states that any Riemannian manifold of dimension $n$ admits an isometric embedding into $\mathbb{R}^{2n+1}$ (Nash, 1954). While it may appear as if this allows for Euclidean embeddings of complex structures with no distortion and only twice the dimension, this is in fact a confusion in vocabulary between the notion of isometric embeddings between those of Riemannian manifolds and those of metric spaces.
@@
In the context of foundational models, we are concerned with globally distance preserving embeddings, as computing the distance between any pairs of token embeddings is crucial for model training. As a result, the Nash Embedding Theorem is not applicable since global distortion could still arise from isometries between Riemannian manifolds. For this reason, by ”isometry”, we refer to those between metric spaces unless otherwise specified, which captures the notion of distortion critical for foundational model embeddings.
@@ A.3 Markov Convexity
In this section we provide the relevant background on notion of Markov convexity. Let $(X,d_{X})$ be a metric space. Then the Markov $p$–convexity constant $\Pi$ (for a fixed positive integer $p$) of the metric space $X$ is a universal constant (or $\infty$) define as follows:
@@ A.4 Non-Euclidean Structure in the Real World
Non-Euclidean Structures in Natural Language Processing. Language exhibits inherently hierarchical structures - from concept taxonomies to entailment relationships - that challenge traditional Euclidean representations. These hierarchical relationships between linguistic units naturally manifest on non-Euclidean manifolds, particularly in hyperbolic space, which has emerged as a powerful framework for natural language processing (Dhingra et al., 2018; Leimeister & Wilson, 2018). Foundational work has demonstrated that hyperbolic embeddings can effectively capture word-level semantics (Tifrea et al., 2019) and concept hierarchies (Le et al., 2019), leveraging the exponential volume growth of hyperbolic space to model tree-like linguistic structures. The success of hyperbolic representations has sparked various advanced applications: from question answering systems (Tay et al., 2018), privacy-preserving text representations (Feyisetan et al., 2019), to multi-document summarization that captures document-level discourse structure (Song et al., 2023). Recent work has further extended these approaches to cross-lingual settings (Saxena et al., 2022) and contextual language models (Chen et al., 2021a), demonstrating the broad utility of non-Euclidean geometries in modern natural language processing.
Non-Euclidean Structures in Computer Vision. Similar to NLP, many computer vision tasks involve data that naturally resides in intricate manifolds that are challenging to model using conventional Euclidean space (Mettes et al., 2024). For instance, visual entities often form inherent hierarchical relationships among object classes, between scenes and their constituent categories (Ge et al., 2023; Pal et al., 2024), or scenes at varying levels of granularity (Khrulkov et al., 2020). In these scenarios, hyperbolic geometry provides a compelling alternative to the Euclidean representations in representing the exponential growth of hierarchical structures with minimal distortion (Sala et al., 2018). Its advantages have been demonstrated across a wide range of applications, including image segmentation (Atigh et al., 2022), action classification (Chen et al., 2022a) video prediction (Surís et al., 2021), deformable 3D surfaces (Masci et al., 2015). In parallel, hyperspherical learning has become integral to modern contrastive learning with cosine similarity, underpining tasks ranging from self-supervised learning (Durrant & Leontidis, 2022) to long-tailed classification (Kasarla et al., 2022) and few-shot learning (Trosten et al., 2023).
@@
Non-Euclidean Structures in Natural Sciences. In natural science, many systems exhibit intricate structures that Euclidean space struggles to capture effectively. In biology, non-Euclidean geometries are integral to analyzing and modeling complex organic structures, such as protein folding (Villegas-Morcillo et al., 2021), single-cell RNA-seq data (Klimovskaia et al., 2020; Ding & Regev, 2021; Bhasker et al., 2024), and phylogenetic trees (Matsumoto et al., 2021), where hyperbolic and spherical geometries are commonly observed. In neuroscience, hyperbolic geometry is shown to be more effective than Euclidean counterpart in modeling the brain’s cortical folding (Urdapilleta et al., 2015), brain surface (Shi et al., 2013), and hippocampal spatial representations (Zhang et al., 2023), aiding in the study of spatial organization and connectivity.
@@ A.5 Deep Learning with Non-Euclidean Geometries
Geodesic Neural Networks. Geodesic neural networks leverage geodesic, particularly geodesic distances, to perform neural operations that preserve geometric structure on manifold-structured data (Bronstein et al., 2017). Several works have developed geodesic convolutional layers by applying filters to local patches in geodesic polar coordinates (Masci et al., 2015), learning directionally sensitive filters along principal curvature directions (Boscaini et al., 2016), or learnable kernel functions that operate on local coordinate systems (Monti et al., 2017). More recent works such as GDGNN (Kong et al., 2022) have integrated geodesic operations with graph representations.
@@
Hyperbolic Neural Networks. Hyperbolic neural networks exploit the geometry of hyperbolic space to learn embeddings that reflect hierarchical relationships more effectively than their Euclidean counterparts (Nickel & Kiela, 2017). HNN (Ganea et al., 2018) and HNN++(Shimizu et al., 2020) developed many basic operations, such as hyperbolic linear and convolutional layers, and multinomial logistic regression (MLR). HGCN (Chami et al., 2019) and HGNN (Liu et al., 2019a) were then among the first to develop hyperbolic graph neural networks (GNNs). More recently, HyboNet (Chen et al., 2021b) proposed a framework of hyperbolic neural networks that does not depend on the Euclidean tangent spaces; Poincaré ResNet (van Spengler et al., 2023) and HCNN (Bdeir et al., 2024) developed components for hyperbolic vision models; LResNet (He et al., 2025) proposed an efficient and stable residual connection method.
@@
Spherical Neural Networks. Spherical neural networks are designed for data that naturally reside on spheres or benefit from spherical symmetry. Spherical CNNs (Cohen et al., 2018; Esteves et al., 2017) extended convolutions and pooling to preserve rotational symmetries. SphereNet (Coors et al., 2018) introduced a framework for learning spherical image representations by encoding distortion invariance into convolutional filters. DeepSphere (Defferrard et al., 2020) proposed a graph-based approach. SWSCNN (Esteves et al., 2020) later proposed a fully spherical CNN that allows for anisotropic filters.
@@
Mix-curvature Neural Networks. Mix-curvature neural networks uses product spaces of the aforementioned manifolds to better model data that have local neighborhoods exhibiting different geometric properties.  (Gu et al., 2019) developed the first learning framework on product spaces, introducing fundamental techniques such as mean and loss functions for embedding optimization. $\kappa$-GCN (Bachmann et al., 2020) then extended learning on product spaces to GCNs, introducing a unified and differentiable Gyrovector spaces framework to constant curvature spaces beyond hyperbolic manifolds.
@@
Non-Euclidean Transformers. Significant advancements have been made toward Transformers in non-Euclidean spaces in recent studies. Within hyperbolic learning, several works have proposed hyperbolic self-attention mechanisms (Gulcehre et al., 2019; Chen et al., 2021b; Shimizu et al., 2020) and hyperbolic linear attentions (Yang et al., 2024c), enabling constructions of hyperbolic Transformers. Hyperbolic fine-tuning methods have also been developed for LLMs (Yang et al., 2024b). Recent works have also proposed hyperbolic vision Transformers (Fein-Ashley et al., 2024). Attention mechanisms have been developed for spherical spaces as well (Lai et al., 2023). Further, Transformers have been developed for mixed curvature manifolds as well (Cho et al., 2023).
@@
Optimization on manifolds. Learning on manifolds often times require optimizing parameters with manifold constraints. Many classical convex optimization algorithms have been extended to the manifold-valued setting (Udriste, 1994; Bacák, 2014; Zhang & Sra, 2016; Weber & Sra, 2022). Stochastic optimization on manifolds has been studied extensivley (Bonnabel, 2013; Zhang et al., 2016; Bécigneul & Ganea, 2018; Weber & Sra, 2021), which includes extensions of algorithms such as SGD and Adam, which are suitable for training models on geometric domains.
@@ Appendix B Additional Statistics and Dataset Details
In this section we give details regarding the datasets we used, as well as the show more statistic results for more LLMs.
B.1 Dataset Details
For the evaluation of token embedding distribution in LLMs, we incorporated a wide range of datasets, including a subset of the RedPajama dataset (Weber et al., 2024) encompassing the arXiv, C4, Common Crawl, GitHub, Wikipedia, and StackExchange datasets; math reasoning datasets such as GSM8K (Cobbe et al., 2021), MATH50K (Hendrycks et al., 2021),MAWPS (Koncel-Kedziorski et al., 2016), and SVAMP (Patel et al., 2021); and common sense reasoning datasets, including BoolQ, WinoGrande (Sakaguchi et al., 2019), and OpenBookQA (Mihaylov et al., 2018).
B.2 More Statistics
In Figure 4 we show the statistics for token embeddings for more LLMs, including GPT-NeoX-20B (Black et al., 2022), OPT-13B (Zhang et al., 2022), RoBERT-Base (Liu et al., 2019b), Gemma2-9B (Team et al., 2024), LLaMa3.1-8B (Grattafiori et al., 2024), and LLaMa-13B (Touvron et al., 2023). The top 2 rows show distribution of the norm of the token embeddings and the bottom 2 rows show the distribution of the frequency of each token embedding. The token frequency distribution demonstrate scale-free property with power law decay, whereas the token norm show rapid decreases in token count for higher normed tokens at the right tail. However, still none of the Euclidean foundational models fully capture the underlying scale-free property of the distribution, with all of them having an initial increase in token count against token norm for small normed token embeddings.
@@ Appendix C $\delta$-Hyperbolicity Computation
Given any four points $a$, $b$, $c$, and $w$ in a metric space, the Gromov product $[a,c]_{w}$ at $w$ is bounded below by the minimum of the Gromov products $[a,b]_{w}$ and $[b,c]_{w}$, minus a slack term $\delta$:
$[a,c]_{w}\geq\min([a,b]_{w},[b,c]_{w})-\delta.$ (1)
@@
$[a,b]_{w}=\frac{1}{2}\left(d(a,w)+d(b,w)-d(a,b)\right).$ (2)
@@
Table 3: Hyperbolicity values $\delta$ for different metric spaces.
Sphere Space Dense Graph PubMed Graph Poincare Space Tree Graph
$\delta$ $0.99\pm 0.01$ $0.62\pm 0.01$ $0.40\pm 0.04$ $0.14\pm 0.01$ $0.0$
@@ Appendix D Foundational Operations for Pretraining Non-Euclidean Foundation Models
Table 4: Geometric Foundation Model Operations: Euclidean vs. Manifold Formulations
@@
  • $\operatorname{PT}_{\mathcal{M}}$:Parallel transport preserving vector properties during translation; $\exp_{\mu_{\mathcal{M}}}$: Exponential map from tangent space at Fréchet mean $\mu_{\mathcal{M}}$; $\log_{\mu_{\mathcal{M}}}$: Inverse exponential map projecting to tangent space; $\operatorname{Ret}$: Retraction mapping for parameter updates; Proj: Tangent space projection operator
Table 4 systematically compares foundational operations in Euclidean space with their adaptations to non-Euclidean manifold spaces, highlighting critical geometric modifications required for pretraining curvature-aware foundation models. Below, we explain the key components and their mathematical formulations:
@@
Optimization. Euclidean SGD $\theta_{t+1}=\theta_{t}-\eta\nabla_{\theta}J(\theta)$ is adapted via retractions $\operatorname{Ret}_{\theta_{t}}$, which map gradient steps $-\eta\text{Proj}_{T_{\theta_{t}}\mathcal{M}}\nabla J$ from the tangent space back to the manifold, ensuring updates respect curvature constraints.
Feed-Forward Network (FFN). Manifold FFNs $y=\exp_{\mathbf{0}}\left(W_{2}\sigma(\log_{\mathbf{0}}(W_{1}\otimes x\oplus b_{1}))\right)$ use Möbius operations ($\otimes$,$\oplus$ , ⊕) for linear transformations and biases, followed by activation in the tangent space. The exponential map $\exp_{\mathbf{0}}$ ensures outputs remain on the manifold.

### v2

@@ A.2 Non-Applicability of the Nash Embedding Theorem
The Nash Embedding Theorem roughly states that any Riemannian manifold of dimension $n$ admits an isometric embedding into $\mathbb{R}^{2n+1}$ [nash1954embedding]. While it may appear as if this allows for Euclidean embeddings of complex structures with no distortion and only twice the dimension, this is in fact a confusion in vocabulary between the notion of isometric embeddings between those of Riemannian manifolds and those of metric spaces.
@@
In the context of foundational models, we are concerned with globally distance preserving embeddings, as computing the distance between any pairs of token embeddings is crucial for model training. As a result, the Nash Embedding Theorem is not applicable since global distortion could still arise from isometries between Riemannian manifolds. For this reason, by "isometry", we refer to those between metric spaces unless otherwise specified, which captures the notion of distortion critical for foundational model embeddings.
@@ A.3 Markov Convexity
In this section we provide the relevant background on notion of Markov convexity. Let $(X,d_{X})$ be a metric space. Then the Markov pp–convexity constant $\Pi$ (for a fixed positive integer $p$) of the metric space $X$ is a universal constant (or $\infty$) define as follows:
@@ A.4 Non-Euclidean Structure in the Real World
Non-Euclidean Structures in Natural Language Processing. Language exhibits inherently hierarchical structures - from concept taxonomies to entailment relationships - that challenge traditional Euclidean representations. These hierarchical relationships between linguistic units naturally manifest on non-Euclidean manifolds, particularly in hyperbolic space, which has emerged as a powerful framework for natural language processing [dhingra2018embedding, leimeister2018skip]. Foundational work has demonstrated that hyperbolic embeddings can effectively capture word-level semantics [tifrea2019poincare] and concept hierarchies [le2019inferring], leveraging the exponential volume growth of hyperbolic space to model tree-like linguistic structures. The success of hyperbolic representations has sparked various advanced applications: from question answering systems [tay2018hyperbolic], privacy-preserving text representations [feyisetan2019leveraging], to multi-document summarization that captures document-level discourse structure [song2023hisum]. Recent work has further extended these approaches to cross-lingual settings [saxena2022cross] and contextual language models [chen2021probing], demonstrating the broad utility of non-Euclidean geometries in modern natural language processing.
Non-Euclidean Structures in Computer Vision. Similar to NLP, many computer vision tasks involve data that naturally resides in intricate manifolds that are challenging to model using conventional Euclidean space [mettes2024hyperbolic]. For instance, visual entities often form inherent hierarchical relationships among object classes, between scenes and their constituent categories [ge2023hyperbolic, pal2024compositional], or scenes at varying levels of granularity [khrulkov2020hyperbolic]. In these scenarios, hyperbolic geometry provides a compelling alternative to the Euclidean representations in representing the exponential growth of hierarchical structures with minimal distortion [sala2018representation]. Its advantages have been demonstrated across a wide range of applications, including image segmentation [atigh2022hyperbolic], action classification [chen2022hmanet] video prediction [suris2021learning], deformable 3D surfaces [masci2015geodesic]. In parallel, hyperspherical learning has become integral to modern contrastive learning with cosine similarity, underpining tasks ranging from self-supervised learning [durrant2022hyperspherically] to long-tailed classification [kasarla2022maximum] and few-shot learning [trosten2023hubs].
@@
Non-Euclidean Structures in Natural Sciences. In natural science, many systems exhibit intricate structures that Euclidean space struggles to capture effectively. In biology, non-Euclidean geometries are integral to analyzing and modeling complex organic structures, such as protein folding [villegas2021foldhsphere], single-cell RNA-seq data [klimovskaia2020poincare, ding2021deep, bhasker2024contrastive], and phylogenetic trees [matsumoto2021novel], where hyperbolic and spherical geometries are commonly observed. In neuroscience, hyperbolic geometry is shown to be more effective than Euclidean counterpart in modeling the brain’s cortical folding [urdapilleta2015can], brain surface [shi2013hyperbolic], and hippocampal spatial representations [zhang2023hippocampal], aiding in the study of spatial organization and connectivity.
@@ A.5 Deep Learning with Non-Euclidean Geometries
Geodesic Neural Networks. Geodesic neural networks leverage geodesic, particularly geodesic distances, to perform neural operations that preserve geometric structure on manifold-structured data [bronstein2017geometric]. Several works have developed geodesic convolutional layers by applying filters to local patches in geodesic polar coordinates [masci2015geodesic], learning directionally sensitive filters along principal curvature directions [boscaini2016learning], or learnable kernel functions that operate on local coordinate systems [monti2017geometric]. More recent works such as GDGNN [kong2022geodesic] have integrated geodesic operations with graph representations.
@@
Hyperbolic Neural Networks. Hyperbolic neural networks exploit the geometry of hyperbolic space to learn embeddings that reflect hierarchical relationships more effectively than their Euclidean counterparts [nickel2017poincare]. HNN [HNN] and HNN++[HNN++] developed many basic operations, such as hyperbolic linear and convolutional layers, and multinomial logistic regression (MLR). HGCN [hgcn2019] and HGNN [liu2019HGNN] were then among the first to develop hyperbolic graph neural networks (GNNs). More recently, HyboNet [chen2021fully] proposed a framework of hyperbolic neural networks that does not depend on the Euclidean tangent spaces; Poincaré ResNet [van2023poincar] and HCNN [Bdeir2024fully] developed components for hyperbolic vision models; LResNet [he2025lresnet] proposed an efficient and stable residual connection method.
@@
Spherical Neural Networks. Spherical neural networks are designed for data that naturally reside on spheres or benefit from spherical symmetry. Spherical CNNs [cohen2018spherical, esteves2017learning] extended convolutions and pooling to preserve rotational symmetries. SphereNet [coors2018spherenet] introduced a framework for learning spherical image representations by encoding distortion invariance into convolutional filters. DeepSphere [deepsphere2020deepsphere] proposed a graph-based approach. SWSCNN [esteves2020swscnn] later proposed a fully spherical CNN that allows for anisotropic filters.
@@
Mix-curvature Neural Networks. Mix-curvature neural networks uses product spaces of the aforementioned manifolds to better model data that have local neighborhoods exhibiting different geometric properties.  [gu2019mixedcurvature] developed the first learning framework on product spaces, introducing fundamental techniques such as mean and loss functions for embedding optimization. $\kappa$-GCN [bachmann2020constant] then extended learning on product spaces to GCNs, introducing a unified and differentiable Gyrovector spaces framework to constant curvature spaces beyond hyperbolic manifolds.
@@
Non-Euclidean Transformers. Significant advancements have been made toward Transformers in non-Euclidean spaces in recent studies. Within hyperbolic learning, several works have proposed hyperbolic self-attention mechanisms [gulcehre2019hyperbolicAT, chen2021fully, HNN++] and hyperbolic linear attentions [yang2024hypformer], enabling constructions of hyperbolic Transformers. Hyperbolic fine-tuning methods have also been developed for LLMs [yang2024hyplora]. Recent works have also proposed hyperbolic vision Transformers [fein-ashley2024hvt]. Attention mechanisms have been developed for spherical spaces as well [lai2023spherical]. Further, Transformers have been developed for mixed curvature manifolds as well [cho2023curve].
@@
Optimization on manifolds. Learning on manifolds often times require optimizing parameters with manifold constraints. Many classical convex optimization algorithms have been extended to the manifold-valued setting [udriste1994convex, bacak2014convex, zhang2016first, weber2022riemannian]. Stochastic optimization on manifolds has been studied extensivley [bonnabel2013stochastic, zhang2016riemannian, becigneul2018riemannian, weber2021projection], which includes extensions of algorithms such as SGD and Adam, which are suitable for training models on geometric domains.
@@ Appendix B Additional Statistics and Dataset Details
In this section we give details regarding the datasets we used, as well as the show more statistic results for more LLMs. We also show the distortion v.s. dimensionality plot for all graph here.
B.1 Distortion v.s. Dimensionality
(a) Tree
(b) Cycle
(c) Ring of Trees
Figure 5: Distortion of embedding a complete tree, cycle, and ring of tree into manifolds of different dimensions (log scale). Each graph has 96 nodes. Euclidean embeddings is shown in blue. In all cases, non-Euclidean geometry achieves significantly smaller distortion with significantly fewer dimensions. The distortion for Euclidean embeddings always plateaus, demonstrating that it is not suited for embeddings each structures regardless of its dimension.
In this section we provide more plots of the distortion of embedding graphs into manifold of varying dimensions. The plots are shown in Figure˜5. In all cases, non-Euclidean geometry achieves significantly smaller distortion with significantly fewer dimensions, reflecting Takeaway 1 in Section˜3.1. The distortion for Euclidean embeddings always plateaus, demonstrating that it is not suited for embeddings each structures regardless of its dimension. On the other hand, the distortion for non-Euclidean embeddings is still being reduced with increased dimensionality for 2 of the structures. This reflects Takeaway 2 in Section˜3.1.
B.2 Dataset Details
For the evaluation of token embedding distribution in LLMs, we incorporated a wide range of datasets, including a subset of the RedPajama dataset [weber2024redpajama] encompassing the arXiv, C4, Common Crawl, GitHub, Wikipedia, and StackExchange datasets; math reasoning datasets such as GSM8K [cobbe2021gsm8k], MATH50K [hendrycksmath2021],MAWPS [koncel2016mawps], and SVAMP [patel2021SVAMP]; and common sense reasoning datasets, including BoolQ, WinoGrande [sakaguchi2019winogrande], and OpenBookQA [mihaylov2018openbookqa].
B.3 More Statistics
In Figure˜6 we show the statistics for token embeddings for more LLMs, including GPT-NeoX-20B [gpt-neox-20b], OPT-13B [zhang2022opt], RoBERT-Base [liu2019roberta], Gemma2-9B [gemma2], LLaMa3.1-8B [dubey2024llama3], and LLaMa-13B [touvron2023llama]. The top 2 rows show distribution of the norm of the token embeddings and the bottom 2 rows show the distribution of the frequency of each token embedding. The token frequency distribution demonstrate scale-free property with power law decay, whereas the token norm show rapid decreases in token count for higher normed tokens at the right tail. However, still none of the Euclidean foundational models fully capture the underlying scale-free property of the distribution, with all of them having an initial increase in token count against token norm for small normed token embeddings.
Table 3: Hyperbolicity values $\delta$ for different metric spaces.
Sphere Space Dense Graph PubMed Graph Poincare Space Tree Graph
$\delta$ $0.99\pm 0.01$ $0.62\pm 0.01$ $0.40\pm 0.04$ $14\pm 0.01 0.0$0.0
@@ Appendix C $\delta$-Hyperbolicity Computation
Given any four points $a$, $b$, $c$, and $w$ in a metric space, the Gromov product $a,c]_{w}$ at $w$ is bounded below by the minimum of the Gromov products $a,b]_{w}$ and $b,c]_{w}$, minus a slack term $\delta$:
$a,c]_{w}\geq\min([a,b]_{w},[b,c]_{w})-\delta.$ (1)
@@
$a,b]_{w}=\frac{1}{2}\left(d(a,w)+d(b,w)-d(a,b)\right).$ (2)
@@ Appendix D Foundational Operations for Pretraining Non-Euclidean Foundation Models
Table 4: Geometric Foundation Model Operations: Euclidean vs. Manifold Formulations. $\operatorname{PT}_{\mathcal{M}}$:Parallel transport preserving vector properties during translation; $\exp_{\mu_{\mathcal{M}}}$: Exponential map from tangent space at Fréchet mean $\mu_{\mathcal{M}}$; $\log_{\mu_{\mathcal{M}}}$: Inverse exponential map projecting to tangent space; $\operatorname{Ret}$: Retraction mapping for parameter updates; Proj: Tangent space projection operator
@@
Table˜4 systematically compares foundational operations in Euclidean space with their adaptations to non-Euclidean manifold spaces, highlighting critical geometric modifications required for pretraining curvature-aware foundation models. Below, we explain the key components and their mathematical formulations:
@@
Optimization. Euclidean SGD $\theta_{t+1}=\theta_{t}-\eta\nabla_{\theta}J(\theta)$ is adapted via retractions $\operatorname{Ret}_{\theta_{t}}$, which map gradient steps $\eta\text{Proj}_{T_{\theta_{t}}\mathcal{M}}\nabla J$ from the tangent space back to the manifold, ensuring updates respect curvature constraints.
Feed-Forward Network (FFN). Manifold FFNs $y=\exp_{\mathbf{0}}\left(W_{2}\sigma(\log_{\mathbf{0}}(W_{1}\otimes x\oplus b_{1}))\right)$ use Möbius operations ($\otimes,\oplus$) for linear transformations and biases, followed by activation in the tangent space. The exponential map $\exp_{\mathbf{0}}$ ensures outputs remain on the manifold.